OpenStudy (anonymous):

1000 female students, with a mean height of 66 inches, and standard deviation of 2. What is the probability of a height between 66 and 70 inches

OpenStudy (anonymous):

You have to use the standard Distribution table with the z-scores for the normal curve. first one is 66 -66/2 = 0.00 second one is 70-66/2 = 2.00 so look on the table fist one is 0.50 second one is 0.9772 just subtract them = 0.9772-0.50 = 0.4772